wiki:UC20200504

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 05/04/2020

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), John Bally (CU), Moire Prescott (NMSU), Mukremin Kilic (OU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Chip Kobulnicky (UWy), Joanne Hughes (Seattle U), Sarah Tuttle (UW), Ben Williams (UW), Derek Buzasi (FGCU)


User feedback and comments from institutional representatives:

  • JHU - nothing to report
  • Washington - nothing to report
  • Colorado - nothing to report
  • UVa - nothing to report
  • FGCU - nothing to report
  • Georgia State - nothing to report
  • Seattle - nothing to report
  • Wyoming - no report
  • NAPG - no report
  • NMSU - nothing to report
  • Oklahoma - nothing to report

Discussion of ARCTIC Bad Columns

Last month Joanne reported some additional bad columns on ARCTIC. Bill and Russet have been investigating this over the past month; Russet summarized their findings. They compared ARCTIC darks from February vs. Joanne’s data in March and identified one new trap column, at the level of about 4 DN. In March the data were acquired in single readout mode whereas the previous data were acquired in quad mode. We know that we get different bad columns/trap columns depending on which amplifier is being used, so could this be the source of the problem? Joanne said no; she switches her observing mode frequently and has not come across this previously. There was also a change in ARCTIC's operating temperature in late February, which could have changed the appearance of bad columns or trap columns. They made some test measurements this weekend, but it looks like the warmer temperature (i.e. what was in place before February) is worse for trap columns and the lower temperature yields fewer counts in the traps (but they don’t go away, but there is also no difference in which columns appeared based on temperature). It could also be due to differences in cosmic rays since they can cause the electronics to misbehave on readout, but this is hard to characterize because CRs are stochastic and there are different readout speeds for different modes. Joanne discovered this issue in an attempt to make a mask to determine when a particular star is on a bad column. The conclusion is that there may be one new trap column but it is likely not related to temperature. We appreciate the feedback from Joanne!


Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 4/02/2020 - 4/28/2020

0) Overview

APO has been closed for all observations due to NM Government restrictions. No observing and absolute minimal staff on site for absolute minimal time. Non-essential systems were powered down or shut off, the telescopes were safed and instruments kept cold and operating except for NICFPS (see below). We are having one person per day visit the site and fill instruments and check the site, and one person making LN2 deliveries weekly. This helps reduce likelihood of catastrophes and provides some basic security, while not increasing the risk of spreading the virus. Note: this is observatory-wide and thus includes SDSS operations as well. The one person who goes to the site each day is being rotated among several of the engineering staff between 3.5m and SDSS. We are expecting the Governor to extend restrictions to May 15, but with the possibility of some easing of restrictions perhaps before then or even extension beyond the 15th. We are making plans to reopen and provide safe operations to staff once we are allowed to do so.

1) Telescope

Nominal operations with minimal problems. Telescope is safed at zenith.

2) Instruments

DIS is powered on and kept cold via cryotiger chillers.

TripleSpec is powered on and cold.

Agile is powered on and ready.

Echelle is powered on and cold.

NIC-FPS is warm and off and will be cooled after we can reopen and replace the ION pump.

ARCTIC is powered on and kept cold via a cryotiger chiller.

ARCSAT: is powered off. Prior to the shut down the new computer was installed in the dome and commissioning started with having two computers splitting functions to improve performance and reduce crashes. We reviewed if we could allow remote operations, however, people felt that this was not recommended for several reasons.


Additional telescope and instrument discussion:

APO is still closed. The NM Governor extended the stay-at-home order to May 15. She eased some restrictions, primarily around retailers. We are still looking at 5/15 as the earliest we could send people back to the site to prepare for operations, and we are currently reviewing a plan of how to do that. The reopening plan will likely include a travel ban for people coming to the site, at least initially. If staff can go back on 5/15, we could maybe get back on the 3.5m that night. We will need to do some engineering early on, so we would aim for resuming science operations that weekend. The one exception is NICFPS: we don’t want to cool it down until we replace the ion pump. It has been ordered and should be in some time this month; this basically constitutes the summer servicing for that instrument.

In response to a query on last month's call we looked at whether we could open up ARCSAT during this period of time, but found that we couldn’t do it under the current guidelines.


APO and the COVID-19 virus

  • Reopening timetable - Mark covered this above. Nancy added that we are working with NMSU, which is developing a phased return to work policy for research labs and off-site research facilities, to make sure our APO reopening plan meets with approval. She also mentioned that this is being addressed across all project directors (3.5m, SDSS IV and V), NMSU, and ARC to ensure everyone is in agreement with our plans. Nancy intends to send an update to the 3.5m user community soon, even if the reopening schedule is still uncertain.
  • Impact on summer shutdown - Mark will contact KPNO this month; if they think we will be rescheduled then we likely won’t plan to realuminize the primary mirror this year.
  • Impact on class trips - We mat not be able to accommodate visitors to the site during Q3 or Q4. Users should notify Nancy, Ben, and Russet if they are planning any class trips to APO for the remainder of the year. We realize that everyone's plans are still uncertain, but the course schedules for the fall semester are already set; if anyone is teaching a class that ordinarily would involve taking students to the site, let us know so we can ensure we have the appropriate remote training materials available. We already know of several planned training visits in June: JHU (a PhD student who will be observing exoplanet transits w/ARCTIC, Kevin observing bright stars with ARCES and an undergraduate helping with both; probably all three will be logged in via TUI from separate locations and videoconferencing with the site staff and each other), and NAPG (Colgate and Bates). We envision that the remote training might be done from three different perspectives: an individual faculty member or postdoc that needs training, a class that will be trained with the ultimate goal of being certified to operate the 3.5m remotely, and a class that would be trained but not for remote observing certification. We would like input from users on what aspects of the training are most important to them.

Q2 3.5m scheduling:

We still have a lot of unassigned time in Q2 (20 slots from end of May to end of June, although a few at the end of June have already been requested). We haven’t been assigning it since we closed down; the hope is that people who lost time may be able to get some back. We would like to receive requests once we know the observatory is reopening (send them to Russet, Ben, Nancy). The institutional schedulers, who should be cc'ed on all requests for unassigned time, should help us determine which are programs are higher priority (e.g. student theses, or more restrictive programs).


Q2 ARCSAT scheduling:

There is one partial unassigned week, otherwise the Q2 schedule is full.


Q3 3.5m scheduling:

Ben is putting together the draft allocations assuming that the summer shutdown will include the primary mirror going out for realuminization. He will get those out later this week so that the institutional schedulers have something to go on, but will indicate that these numbers are lower limits and are subject to change. Requests will be due at the end of May. Proposers should indicate if they have any flexibility in the scheduling of their programs (e.g. would like dark time but could use grey).


KOSMOS Update:

No progress has been made because the TEG cannot access their labs. During a phased return to research facilities they can practice socially distant diagnostics of the dewar, detector, and software.


ARCTIC Diffuser Rotation

We had a Preliminary Design Review a week ago, which went well. Kal is working to address a few issues that were raised during the PDR and is getting additional quotes for having some parts made, but we are cautiously optimistic that the diffuser will be rotating again in the near future. The TEG members can go to work to retrieve things, so it's possible that they could do some home assembly if need be.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: the DAWG met recently but Nancy did not get an update. The group seems to be compiling a comprehensive list of quick-look and more extensive data analysis pipelines for various instruments. They also intend to start writing the white paper. Nancy is hoping they will also be considering the issue of data storage, backup, archiving etc. on the observatory side so that she can use that input to inform decisions about future spending on IT infrastructure.
  • Getting narrow-band filters for ARCTIC: Still in Nancy's court; she will try again to contact the vendor in Japan now that the pandemic is less severe there.
  • Echelle Advisory Group: a group of people met to brainstorm ideas for a new instrument that could ultimately replace ARCES. Several leads are still being investigated but users are asked to let Nancy know if there are new science drivers that should be taken into consideration.

ACTION ITEMS:

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN.

New action items from this meeting:

  • Users should notify Nancy, Ben, and Russet if they are planning any class trips to APO for the remainder of the year.
  • Users are asked to provide input on what aspects of the on-site training are most important to them so that the remote training materials can be prioritized.
  • Proposers for Q3 time are asked indicate the degree of flexibility they have in the scheduling of their requested observations.

All Other Business

None.


Next meeting:

The next Users Committee Zoom meeting will be on Tuesday June 2 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.